# On the kernels of the pro-$p$ outer Galois representations associated to once-punctured CM elliptic curves

### v2

@@ 3.2. Two-variable filtration on pro-$p$ mapping class group
Here, we use $f(\alpha)\alpha^{-1}\in\Pi(\boldsymbol{m}+\boldsymbol{1})$, which follows from Lemma 3.5, to deduce the last congruence. By Lemma 3.8, the last term is equal to $\chi_{1}^{m_{1}}(\gamma)(f(y)y^{-1})$ as desired.
Secondly, we compute the term $(\gamma\cdot f)(x)x^{-1}$ as
@@
By Lemma 3.7, the element $\beta\coloneqq x^{-\chi_{1}(\gamma^{-1})}\gamma^{-1}(x)$ is contained in the subgroup $\Pi(2,0)$. Since $f(\beta)\beta^{-1}$ is contained in $\Pi(\boldsymbol{m}+(2,0))$ by Lemma 3.5, we have
@@
for every $n\in\mathbb{Z}_{p}$. By continuity, it suffices to prove the congruence for every $n\in\mathbb{Z}_{\geq 1}$, which we prove by induction on $n$. We can compute the term $f(x^{n})x^{-n}$ as follows:
@@
$[f(x),(f(x)x^{-1})^{n-1}]\in\Pi(\boldsymbol{m}+(2,0))$
to establish the last congruence. ∎
Remark 3.10.
@@
By Proposition 3.9, we now understand how the group $\Gamma^{{\dagger}}_{1}/F^{(1,0)}\Gamma^{{\dagger}}$ acts on $gr^{\boldsymbol{m}}_{1}\tilde{\Gamma}$. The following lemma allows us to describe the action of $\Gamma_{1}/F^{(1,0)}\Gamma$ on $gr^{\boldsymbol{m}}_{1}\Gamma$:
@@
The natural homomorphism
$\Gamma^{{\dagger}}_{1}/F^{(1,0)}\Gamma^{{\dagger}}\to\Gamma_{1}/F^{(1,0)}\Gamma$
is an isomorphism. Similarly, we have $\Gamma^{{\dagger}}_{2}/F^{(0,1)}\Gamma^{{\dagger}}\xrightarrow{\sim}\Gamma_{2}/F^{(0,1)}\Gamma$.
@@
Since $\ker(\Gamma^{{\dagger}}_{1}\to\Gamma_{1})=\ker(F^{(1,0)}\Gamma^{{\dagger}}\to F^{(1,0)}\Gamma)=\langle\mathrm{inn}(z)\rangle$, it suffices to prove that both $\Gamma^{{\dagger}}_{1}\to\Gamma_{1}$ and $F^{(1,0)}\Gamma^{{\dagger}}\to F^{(1,0)}\Gamma$ are both surjective by the snake lemma. First, we prove the surjective of $\Gamma^{{\dagger}}_{1}\to\Gamma_{1}$. Let $\bar{\gamma}\in\Gamma_{1}$ be an arbitrary element and fix an arbitrary lift $\gamma\in\tilde{\Gamma}_{1}$ of $\bar{\gamma}$. If we write $\gamma(z)=gz^{\alpha}g^{-1}$ for some $g\in\Pi$ and $\alpha\in\mathbb{Z}_{p}^{\times}$, $g^{-1}\gamma g=\mathrm{inn}(g^{-1})\circ\gamma$ preserves $\langle z\rangle$. Moreover,
@@
This shows $g^{-1}\gamma g\in\Gamma^{{\dagger}}_{1}$, hence the surjectivity of $\Gamma^{{\dagger}}_{1}\to\Gamma_{1}$.
As for the surjectivity of $F^{(1,0)}\Gamma^{{\dagger}}\to F^{(1,0)}\Gamma$, let $\bar{\gamma}\in F^{(1,0)}\Gamma$ be arbitrary element and fix an arbitrary lift $\gamma\in F^{(1,0)}\tilde{\Gamma}$ of $\bar{\gamma}$. If we write $\gamma(z)=gz^{\alpha}g^{-1}$ for some $g\in\Pi$ and $\alpha\in\mathbb{Z}_{p}^{\times}$, then $g^{-1}\gamma g\in\Gamma^{{\dagger}}_{1}$ by the above argument. Observe that
@@
Hence, to prove $(g^{-1}\gamma g)(x)x^{-1}=[g^{-1},\gamma(x)]\gamma(x)x^{-1}\in\Pi(2,0)$, it is enough to show $g\in\Pi(1,0)$. Note that $\Pi(\boldsymbol{1})/\Pi(2,1)=\Pi(\boldsymbol{1})/\Pi(2,0)\subset\Pi(1,0)/\Pi(2,0)$ is a free $\mathbb{Z}_{p}[[\Pi/\Pi(1,0)]]$-submodule generated by $z=(y-1)x$. Hence $\gamma(z)z^{-1}=[g,z]=(g-1)z\in\Pi(2,0)$ implies $g=1$ in $\mathbb{Z}_{p}[[\Pi/\Pi(1,0)]]$, i.e. $g\in\Pi(1,0)$. This concludes the proof. ∎
@@
For every $\boldsymbol{m}=(m_{1},m_{2})\in\mathbb{Z}_{\geq 0}^{2}\setminus\{(0,0)\}$, the action of $\Gamma_{1}/F^{(1,0)}\Gamma$ (resp. $\Gamma_{2}/F^{(0,1)}\Gamma$) on $\mathrm{gr}^{\boldsymbol{m}}_{1}\Gamma$ (resp. $\mathrm{gr}^{\boldsymbol{m}}_{2}\Gamma$) is given by $\chi_{1}^{m_{1}}$ (resp. $\chi_{2}^{m_{2}}$).
@@
$F^{(1,0)}\Gamma^{{\dagger}}=\ker(\Gamma^{{\dagger}}_{1}\xrightarrow{\chi_{1}}\mathbb{Z}_{p}^{\times})\quad\text{and}\quad F^{(0,1)}\Gamma^{{\dagger}}=\ker(\Gamma^{{\dagger}}_{2}\xrightarrow{\chi_{2}}\mathbb{Z}_{p}^{\times}).$
@@
We only prove the first equality. Let $\gamma$ be an element of $\ker(\Gamma^{{\dagger}}_{1}\xrightarrow{\chi_{1}}\mathbb{Z}_{p}^{\times})$. Since we have $\gamma(y)y^{-1}\in\Pi(2)=\Pi(1,1)$ by definition of $\Gamma^{{\dagger}}_{1}$, it suffices to show $\gamma(x)=x\bmod\Pi(2,0)$.
Since the action of $\Gamma^{{\dagger}}_{1}$ on $\Pi(1,0)/\Pi(2,0)$, which is a free $\mathbb{Z}_{p}[[\Pi/\Pi(1,0)]]$-module generated by the image of $x$ is $\mathbb{Z}_{p}[[\Pi/\Pi(1,0)]]$-linear, the desired equality follows from the equality $\gamma(z)=z$. ∎
@@ 3.3. Two-variable filtration on Galois group
$\rho_{X,p}\colon G_{K}\to\mathrm{Out}(\Pi_{1,1}),$
and the image of $\rho_{X,p}$ is contained in $\Gamma\subset\Gamma_{1,1}$ introduced in the last subsection.
For $\boldsymbol{m}\in\mathbb{Z}_{\geq 0}^{2}\setminus\{(0,0)\}$, let $F^{\boldsymbol{m}}G_{K}\subset F^{\lvert\boldsymbol{m}\rvert}G_{K}$ denote the inverse image of $F^{\boldsymbol{m}}\Gamma$ under $\rho_{X,p}$. We also define the associated graded quotients by
$\mathrm{gr}^{\boldsymbol{m}}_{1}G_{K}\coloneqq F^{\boldsymbol{m}}G_{K}/F^{\boldsymbol{m}+(1,0)}G_{K}\quad\text{and}\quad\mathrm{gr}^{\boldsymbol{m}}_{2}G_{K}\coloneqq F^{\boldsymbol{m}}G_{K}/F^{\boldsymbol{m}+(0,1)}G_{K}.$
Moreover, let $F_{1}G_{K}$ (resp. $F_{2}G_{K}$) denote the inverse image of $\Gamma_{1}$ (resp. $\Gamma_{2}$).
@@
The assertion of (1) is clear. By Lemma 3.13 and Lemma 3.11, the group $F^{(1,0)}G_{K}$ coincides with the kernel of $F_{1}G_{K}\xrightarrow{\chi_{1}}\mathbb{Z}_{p}^{\times}$, and thus it is equal to $G_{K(E[p^{\infty}])}$. The same argument shows $F^{(0,1)}G_{K}=G_{K(E[p^{\infty}])}$. Since we have
@@
the last assertion follows. ∎
@@
For every $\boldsymbol{m}\in\mathbb{Z}_{\geq 0}^{2}\setminus\{(0,0)\}$, the action of $\mathrm{Gal}(K(E[p^{\infty}])/K(E[\bar{\mathfrak{p}}^{\infty}]))$ on $\mathrm{gr}^{\boldsymbol{m}}_{1}G_{K}$ is given by the character $\chi_{1}^{m_{1}}$. Similarly, the action of $\mathrm{Gal}(K(E[p^{\infty}])/K(E[\mathfrak{p}^{\infty}]))$ on $\mathrm{gr}^{\boldsymbol{m}}_{2}G_{K}$ is given by $\chi_{2}^{m_{2}}$.
@@
4. Proof of main theorem
In this section, we assume that $p$ satisfies the first two assumptions of Theorem 2.12, i.e. the class number of $K(p)$ is not divisible by $p$ and there are exactly two primes of $K(p^{\infty})$ above $p$. Moreover, we abbreviate $\mathrm{Gal}(\Omega/K(p))$ and $\mathrm{Gal}(\Omega^{\ast}/K(p))$ as $G$ and $G^{\ast}$, respectively.
Let us recall the definition of the index set
@@ 4. Proof of main theorem
In this section, we also use the following subset of $I$:
@@
First, we give two filtrations on $G^{\ast}$. By Lemma 2.11, we have
$\rho_{X,p}(G_{K(p)})=\mathrm{Gal}(\Omega^{\ast}K(E[p])/K(p))=G^{\ast}\times\mathrm{Gal}(K(E[p])/K(p))\subset\Gamma_{1,1}.$
@@
Moreover, by pulling back these filtrations under the projection $G\to G^{\ast}$, we obtain two filtrations $\{F^{m}G\}_{m\geq 1}$ and $\{F^{\boldsymbol{m}}G\}_{\boldsymbol{m}}$ on $G$. Then, by construction, the graded Lie algebras associated to $\{F^{m}G^{\ast}\}_{m\geq 1}$ and $\{F^{m}G\}_{m\geq 1}$ are naturally isomorphic to $\mathfrak{g}=\oplus_{m\geq 1}\mathfrak{g}_{m}$ defined in Section 2.2.
In the subsequent subsections, we give a proof of Theorem 2.12. The proof follows Sharifi’s approach [Sha02, Theorem 1.2]: We show that the intersection $\cap_{m\geq 1}F^{m}G=\mathrm{Gal}(\Omega/\Omega^{\ast})$ is trivial by the following strategy.
@@
    In Section 4.1, we construct an element $\sigma_{\boldsymbol{m}}\in F^{\boldsymbol{m}}G$ whose image in $\mathfrak{g}_{|\boldsymbol{m}|}$ satisfies the assumption of Conjecture 2.8. Moreover, from its construction, it is proved in Section 4.2 that $\{\sigma_{\boldsymbol{m}}\}_{\boldsymbol{m}\in I}$ strongly generates $F^{1}G$. Here, properties of two-variable filtrations are used in an essential way.
@@
    Conjecture 2.8 implies that the filtration $\{F^{m}G\}_{m\geq 1}$ coincides with the ‘‘fastest’’ descending central filtration $\{\tilde{F}^{m}G\}_{m\geq 1}$ satisfying $\sigma_{\boldsymbol{m}}\in\tilde{F}^{|\boldsymbol{m}|}G$ for every $\boldsymbol{m}\in I$. Since the intersection of the latter filtration is proved to be trivial in Lemma 4.11 (3), we obtain the desired result.

### v3

@@ 3.2. Two-variable filtration on pro-$p$ mapping class group
Here, we use $f(\alpha)\alpha^{-1}\in\Pi(\boldsymbol{m}+\boldsymbol{1})$, which follows from Lemma 3.9, to deduce the last congruence. By Lemma 3.13, the last term coincides with $\chi_{1}^{m_{1}}(\gamma)(f(y)y^{-1})$ as desired. Secondly, we compute $(\gamma\cdot f)(x)x^{-1}$ as
@@
By Lemma 3.12, the element $\beta\coloneqq x^{-\chi_{1}(\gamma^{-1})}\gamma^{-1}(x)$ is contained in $\Pi(2,0)$. Since $f(\beta)\beta^{-1}$ is contained in $\Pi(\boldsymbol{m}+(2,0))$ by Lemma 3.9, we have
@@
for every $n\in\mathbb{Z}_{p}$. By continuity, it suffices to prove the congruence for every $n\in\mathbb{Z}_{\geq 1}$, which we prove by induction on $n$. We compute $f(x^{n})x^{-n}$ as follows:
@@
$f(x),(f(x)x^{-1})^{n-1}]\in\Pi(\boldsymbol{m}+(2,0))$
to establish the last congruence. This concludes the proof. ∎
Remark 3.14.1.
@@
By Proposition 3.14, we now understand how $\Gamma^{{\dagger}}_{1}/F^{(1,0)}\Gamma^{{\dagger}}$ acts on $gr^{\boldsymbol{m}}_{1}\tilde{\Gamma}$. The following lemma allows us to describe the action of $\Gamma_{1}/F^{(1,0)}\Gamma$ on $gr^{\boldsymbol{m}}_{1}\Gamma$:
@@
The natural homomorphisms
$\Gamma^{{\dagger}}_{1}/F^{(1,0)}\Gamma^{{\dagger}}\to\Gamma_{1}/F^{(1,0)}\Gamma\quad\text{and}\quad\Gamma^{{\dagger}}_{2}/F^{(0,1)}\Gamma^{{\dagger}}\to\Gamma_{2}/F^{(0,1)}\Gamma$
are isomorphisms.
@@
Since we have
$\operatorname{Ker}(\Gamma^{{\dagger}}_{1}\to\Gamma_{1})=\operatorname{Ker}(F^{(1,0)}\Gamma^{{\dagger}}\to F^{(1,0)}\Gamma)=\langle\mathrm{inn}(z)\rangle,$
it suffices to prove that both $\Gamma^{{\dagger}}_{1}\to\Gamma_{1}$ and $F^{(1,0)}\Gamma^{{\dagger}}\to F^{(1,0)}\Gamma$ are surjective. First, we prove the surjectivity of $\Gamma^{{\dagger}}_{1}\to\Gamma_{1}$. Let $\bar{\gamma}\in\Gamma_{1}$ be an arbitrary element and fix an arbitrary lift $\gamma\in\tilde{\Gamma}_{1}$ of $\bar{\gamma}$. If we write $\gamma(z)=gz^{\alpha}g^{-1}$ for some $g\in\Pi$ and $\alpha\in\mathbb{Z}_{p}^{\times}$, $g^{-1}\gamma g=\mathrm{inn}(g^{-1})\circ\gamma$ preserves $\langle z\rangle$. Moreover,
@@
This shows $g^{-1}\gamma g\in\Gamma^{{\dagger}}_{1}$, hence the surjectivity of $\Gamma^{{\dagger}}_{1}\to\Gamma_{1}$. Regarding the surjectivity of $F^{(1,0)}\Gamma^{{\dagger}}\to F^{(1,0)}\Gamma$, let $\bar{\gamma}\in F^{(1,0)}\Gamma$ be arbitrary element and fix a lift $\gamma\in F^{(1,0)}\tilde{\Gamma}$ of $\bar{\gamma}$. If we write $\gamma(z)=gz^{\alpha}g^{-1}$ for some $g\in\Pi$ and $\alpha\in\mathbb{Z}_{p}^{\times}$, then we have $g^{-1}\gamma g\in\Gamma^{{\dagger}}_{1}$ by the above argument. Observe that
@@
Hence, to prove $(g^{-1}\gamma g)(x)x^{-1}=[g^{-1},\gamma(x)]\gamma(x)x^{-1}\in\Pi(2,0)$, it is enough to show $g\in\Pi(1,0)$. Note that $\Pi(\boldsymbol{1})/\Pi(2,1)=\Pi(\boldsymbol{1})/\Pi(2,0)\subset\Pi(1,0)/\Pi(2,0)$ is a free $\mathbb{Z}_{p}[[\Pi/\Pi(1,0)]]$-submodule generated by z=(y1)xz=(y-1)x. Hence that $\gamma(z)z^{-1}=[g,z]=(g-1)z$ is contained in $\Pi(2,0)$ implies $g=1$ in $\mathbb{Z}_{p}[[\Pi/\Pi(1,0)]]$, i.e. $g\in\Pi(1,0)$. This concludes the proof. ∎
@@
Let $\boldsymbol{m}=(m_{1},m_{2})\in\mathbb{Z}_{\geq 0}^{2}\setminus\{(0,0)\}$. The action of $\Gamma_{1}/F^{(1,0)}\Gamma$ (resp. $\Gamma_{2}/F^{(0,1)}\Gamma$) on $\operatorname{Gr}^{\boldsymbol{m}}_{1}\Gamma$ (resp. $\operatorname{Gr}^{\boldsymbol{m}}_{2}\Gamma$) is given by $\chi_{1}^{m_{1}}$ (resp. $\chi_{2}^{m_{2}}$).
@@
$F^{(1,0)}\Gamma^{{\dagger}}=\operatorname{Ker}(\Gamma^{{\dagger}}_{1}\xrightarrow{\chi_{1}}\mathbb{Z}_{p}^{\times})\quad\text{and}\quad F^{(0,1)}\Gamma^{{\dagger}}=\operatorname{Ker}(\Gamma^{{\dagger}}_{2}\xrightarrow{\chi_{2}}\mathbb{Z}_{p}^{\times}).$
@@
We only prove the first equality. Let $\gamma$ be an element of $\operatorname{Ker}(\Gamma^{{\dagger}}_{1}\xrightarrow{\chi_{1}}\mathbb{Z}_{p}^{\times})$. Since we have $\gamma(y)y^{-1}\in\Pi(2)=\Pi(1,1)$, it suffices to show $\gamma(x)=x\bmod\Pi(2,0)$. Since the action of $\Gamma^{{\dagger}}_{1}$ on $\Pi(1,0)/\Pi(2,0)$, which is a free $\mathbb{Z}_{p}[[\Pi/\Pi(1,0)]]$-module generated by the image of $x$, is $\mathbb{Z}_{p}[[\Pi/\Pi(1,0)]]$-linear, the desired equality follows from the equality $\gamma(z)=z$. ∎
@@ 3.3. Two-variable filtration on Galois group
$\rho_{X,p}\colon G_{K}\to\operatorname{Out}(\Pi_{1,1}),$
and the image of $\rho_{X,p}$ is contained in $\Gamma\subset\Gamma_{1,1}$ introduced in the last subsection. For $\boldsymbol{m}\in\mathbb{Z}_{\geq 0}^{2}\setminus\{(0,0)\}$, let $F^{\boldsymbol{m}}G_{K}\subset F^{\lvert\boldsymbol{m}\rvert}G_{K}$ denote the inverse image of $F^{\boldsymbol{m}}\Gamma$ under $\rho_{X,p}$. We define the associated graded quotients by
$\operatorname{Gr}^{\boldsymbol{m}}_{1}G_{K}\coloneqq F^{\boldsymbol{m}}G_{K}/F^{\boldsymbol{m}+(1,0)}G_{K}\quad\text{and}\quad\operatorname{Gr}^{\boldsymbol{m}}_{2}G_{K}\coloneqq F^{\boldsymbol{m}}G_{K}/F^{\boldsymbol{m}+(0,1)}G_{K}.$
Moreover, let $F_{1}G_{K}$ (resp. $F_{2}G_{K}$) be the inverse image of $\Gamma_{1}$ (resp. $\Gamma_{2}$).
@@
The assertion of (1) is clear. By Lemma 3.15 and Lemma 3.17, the group $F^{(1,0)}G_{K}$ coincides with the kernel of $F_{1}G_{K}\xrightarrow{\chi_{1}}\mathbb{Z}_{p}^{\times}$, and thus it coincides with $G_{K(E[p^{\infty}])}$. The same argument shows $F^{(0,1)}G_{K}=G_{K(E[p^{\infty}])}$. Since we have
@@
the last assertion also follows. ∎
The following corollary is one of key ingredients to prove Theorem 2.14:
@@
Let $\boldsymbol{m}\in\mathbb{Z}_{\geq 0}^{2}\setminus\{(0,0)\}$. The action of $\operatorname{Gal}(K(E[p^{\infty}])/K(E[\bar{\mathfrak{p}}^{\infty}]))$ on $\operatorname{Gr}^{\boldsymbol{m}}_{1}G_{K}$ is given by the character $\chi_{1}^{m_{1}}$. Similarly, the action of $\operatorname{Gal}(K(E[p^{\infty}])/K(E[\mathfrak{p}^{\infty}]))$ on $\operatorname{Gr}^{\boldsymbol{m}}_{2}G_{K}$ is given by $\chi_{2}^{m_{2}}$.
@@
4. Proof of main result
In this section, we keep the same notation as in Section 2. Moreover, we abbreviate $\operatorname{Gal}(\Omega/K(p))$ and $\operatorname{Gal}(\Omega^{\ast}/K(p))$ as $G$ and $G^{\ast}$, respectively. Recall that the index set $I$ is defined by
@@ 4. Proof of main result
In this section, we also use the following subset:
@@
We first give two filtrations on $G^{\ast}$. By Lemma 2.13, we have
$\rho_{X,p}(G_{K(p)})=G^{\ast}\times\operatorname{Gal}(K(E[p])/K(p))\subset\Gamma_{1,1}.$
@@
Moreover, by taking pullbacks of these filtrations under the natural projection, we obtain filtrations $\{F^{m}G\}_{m\geq 1}$ and $\{F^{\boldsymbol{m}}G\}_{\boldsymbol{m}}$. Then, by construction, the graded Lie algebras associated to $\{F^{m}G^{\ast}\}_{m\geq 1}$ and $\{F^{m}G\}_{m\geq 1}$ are isomorphic to $\mathfrak{g}=\oplus_{m\geq 1}\mathfrak{g}_{m}$ defined in Section 2.1.
In the subsequent subsections, we give a proof of Theorem 2.14. The proof follows Sharifi’s approach [Sha02, Theorem 1.2]. Namely, we show that the intersection $\cap_{m\geq 1}F^{m}G=\operatorname{Gal}(\Omega/\Omega^{\ast})$ is trivial by the following strategy.
@@
    In Section 4.1, we construct an element $\sigma_{\boldsymbol{m}}\in F^{\boldsymbol{m}}G$ whose image in $\mathfrak{g}_{|\boldsymbol{m}|}$ satisfies the assumption of Conjecture 2.3. Moreover, it is proved in Section 4.2 that $\{\sigma_{\boldsymbol{m}}\}_{\boldsymbol{m}\in I}$ strongly generates $F^{1}G$. Here, several properties of two-variable filtrations are used in an essential way.
@@
    Conjecture 2.3 implies that $\{F^{m}G\}_{m\geq 1}$ coincides with the ‘‘fastest’’ descending central filtration $\{\tilde{F}^{m}G\}_{m\geq 1}$ satisfying $\sigma_{\boldsymbol{m}}\in\tilde{F}^{|\boldsymbol{m}|}G$ for every $\boldsymbol{m}\in I$. Since the intersection of the latter filtration is proved to be trivial by Lemma 4.12 (3) in a purely group-theoretic way, we obtain the desired result.
We first prepare a lemma. This allows us to apply Proposition 2.11 to construct a basis of $\mathfrak{g}\otimes\mathbb{Q}_{p}$ under the assumption of Theorem 2.14:
Lemma 4.1.
Assume that the class number of $K(p)$ is not divisible by $p$ and there are exactly two primes of $K(p^{2})$ above $p$. Then the character $\kappa_{\boldsymbol{m}}$ is nontrivial for every $\boldsymbol{m}\in I$.
Proof.
The assertion follows from Theorem 2.9 and Corollary A.2. ∎